12) A researcher measures the amount of sugar in several cans of the same soda. The mean is 39.01 with a standard deviation of 0.5. The researcher randomly selects a sample of 100.

Find the sum with a z-score of −2.4. (Enter an exact number as an integer, fraction, or decimal.)
Answer #1

Solution :

Given that ,

mean = = 39.01

standard deviation = = 0.5

n = 100

= 39.01

=  / n = 0.5/ 100=0.05

= z =-2.4

= z * +

= -2.4*0.05+39.01

= 38.89